OpenStudy (anonymous):

For the line l passing through the points (1,0,1) and (2,4,7) what is the value of c in the following symmetric equations for l? I need help I could not solve it

OpenStudy (anonymous):

x-1=y/4=z-1\c

OpenStudy (anonymous):

wat is meaning of symetric here .. are they parrellel

OpenStudy (anonymous):

yes they are

OpenStudy (anonymous):

d.r's are <x 2-x1,y2-y1,z2-z1> i.e., <2-1,4-0,7-1> or < 1,4,6 > eq. of line is \[\frac{ x-1 }{ 1 }=\frac{ y-0 }{ 4 }=\frac{ z-1 }{ 6 }\]